# Dynamics in an emergent quantum-like state space generated by a nonlinear classical network

### v3

@@ IV.1 Phase and the adjacency matrix
This unitary transformation of $A$ does not change the spectrum, but it does rotate the basis for the eigenstates. We use this transformation below, after defining projection of the emergent states, to show how the eigenvectors of the projected emergent states transform under continuous rotations. Later in this paper, we use the transformation of Eq. 13 to account for how oscillator phases in the classical network dictate the eigenvector basis. That basis change, in general, varies throughout the ensemble and therefore influences the mixedness of the state[40]. Recall that we use the term ‘ensemble’ to mean an average over initial phases to account for state preparation.
@@
In the numerical examples shown below we obtain the phases from simulations using the Kuramoto model, described in the previous section. For any time point during the simulation, we use the oscillator phases to define $\Phi$, then we diagonalize $\Phi^{-1}A\Phi$ to obtain the eigenvectors of the emergent states—now with complex coefficients. We project these coefficients to those for an effective two-state system for each QL bit, as we now describe for the example of a single QL bit. The method is generalized in the obvious way for a system comprising any number of QL bits, as shown in Sec. V.C.
IV.2 The effective 2-state model for QL bit emergent states
@@ IV.3 Phase in the states
For a synchronized oscillator system on the QL bit graph with all edges set to 1 in the adjacency matrix, we find that $\alpha_{0}=\beta_{0}\approx 1/\sqrt{2}$ and $\alpha_{1}=-\beta_{1}\approx 1/\sqrt{2}$ (we use the convention that $\beta_{1}$ takes the negative sign). More generally, we can bias the edges that connect the two subgraphs by giving them values (in the adjacency matrix) of any value on the unit circle in the complex plane. Recall that the special case of negative-signed edges connecting the subgraphs can be interpreted as phases of the two subgraphs being perfectly out of phase[1]. The resulting eigenvaectors can be evaluated by considering the effective adjacency matrix for the QL bit graph comprising two $d$-regular subgraphs coupled by edges given the bias $e^{i\alpha}$:
@@
By rotating the edge bias, we therefore generate states (when the system is perfectly synchronized) that emulate, for example, the states of a system of spins or polarization states of photons. The notable emergent states isomorphic to polarization states of light are: (i) the linear combinations $a_{1}\pm a_{2}$, when edges take values ±$\pm 1$ respectively, which are comparable to linear combinations of vertical and horizontal polarization; (ii) the states $a_{1}\pm ia_{2}$, when edges take values of ±$\pm i$, analogous to right- and left-circularly polarized light.
IV.4 Synchronization of quantum bits
@@
with the $\mathbf{L}$ being the $4\times 4$ skew symmetric matrices[45]:
@@
V Synchronization versus decoherence of QL states
V.1 Methods
Here we study a phase oscillator network model with a coupling structure templated by the Cartesian product of two QL bit graphs $G_{A}\Box G_{B}$. The phase oscillators are indicated by vertices of the graph, and are assigned a phase that biases the edges, as we describe below.
@@ V.1 Methods
V.2 Parameters and State preparation
@@
In more detail, whether or not a system of oscillators will perfectly synchronize depends not only on $K$ and $\sigma_{\nu}$, but also on the initial conditions. That is, the specific assignment of frequencies and initial phases to each oscillator in the network. For insight into this statement see, for example, Refs. [10, 11, 9, 47, 48]. This sensitivity of the system to initial conditions emphasizes the nonlinearity of the network.
V.3 Observation of a State
According to Eq. 12 we are applying a unitary transformation to the adjacency matrix that is decided by the oscillators phases at any time. Therefore the spectrum does not change with order parameter, only the state. The density matrix for the relevant emergent state is calculated in the basis produced by the effective two-state model for each QL bit. That is, we statistically average over the pure states:
$W_{i}^{4\times 4}(t)=c^{i}_{1}(t)|a_{1}\rangle|b_{1}\rangle+c^{i}_{2}(t)|a_{2}\rangle|b_{1}\rangle+c^{i}_{3}(t)|a_{1}\rangle|b_{2}\rangle c^{i}_{4}(t)|a_{2}\rangle|b_{2}\rangle,$ (25)
@@ V.3 Observation of a State
for each time point in a simulation. Here the $c_{k}(t)$ are complex coefficients obtained from the projections analogous to those outlined in Sec. IV.B. We obtain the coefficients as follows. We form the Cartesian product of two QL bit graphs, say $G_{A}$ and $G_{B}$, each comprising two subgraphs with $n_{0}$ vertices. From the product, $G_{A}\Box G_{B}$, we end up with four subgraphs (each with $n_{0}^{2}$ vertices). The subgraphs are connected with a certain structure of edges that define the correlations. The vertices in each subgraph are labeled by all possible $n_{0}^{2}$ pairs of vertex indices, where one index is taken from a subgraph of $G_{A}$ and the other from a subgraph of $G_{B}$. We project the emergent states of the entire graph to effective states in a chosen basis. Here we choose the natural product basis associated directly with the subgraphs of $G_{A}\Box G_{B}$: $|a_{1}\rangle|b_{1}\rangle$, $|a_{2}\rangle|b_{1}\rangle$, $|a_{1}\rangle|b_{2}\rangle$, $|a_{2}\rangle|b_{2}\rangle$. That basis is selected by defining the vectors
@@
where $\sum_{\text{prep}}$ means an average over state preparations. We report $\mathrm{Tr}(\rho)^{2}$ as a simple measure of mixedness of the state. Notice that the complex phases cancel to unity in the diagonal elements of $\rho_{mn}(t)$. On the other hand, the complex phases are retained in the off-diagonal elements and unless the system is synchronized, they interfere in the average and can thereby diminish or erase the off-diagonal elements, introducing state mixedness.
The vectors $J_{a1b1}$ and so on define the measurement basis for the QL state. Here we set them up so that the basis is the product basis associated directly with the subgraphs of $G_{A}\Box G_{B}$. We could, however, choose any other basis. For instance, we could select the Bell basis using
$J_{a1b1\pm a2b2}=\{\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\underbrace{0,0,\dots,0}_{2n_{0}^{2}\text{ times}}\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\}/\sqrt{N}$ (33)
$J_{a1b2\pm a2b1}=\{\underbrace{0,0,\dots,0}_{n_{0}^{2}\text{ times}}\underbrace{\pm 1,\pm 1,\dots,\pm 1}_{n_{0}^{2}\text{ times}}\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\underbrace{0,0,\dots,0}_{n_{0}^{2}\text{ times}}\}/\sqrt{N}.$ (34)
A preferred basis could be set up by suitable weighting of the edges.
V.4 Synchronization
@@
Despite the fact that the underlying synchronization dynamics are inherently nonlinear, maps in the state space taking the state from one time point to another, $\Xi:\rho(t_{1})\rightarrow\rho(t_{2})$, are linear maps.
@@
Moreover, the $\Xi$ are trace-preserving completely positive maps[50, 51, 52, 53] because they map $\rho\rightarrow\rho^{\prime}$. However, the dynamical evolution is generally not unitary; being unitary only in the limit of strong synchronization.
Different regimes can be identified. When all the oscillators are strongly synchronized within each QL bit, but not necessarily between QL bits, then the general model for the single-excitation subspace specializes to the Lohe model. Single-excitation subspace means that the network graph carries the structure of the tensor sum of QL bit graphs and QL bits are coupled to other QL bits. When the coupling $K$ for the classical system is very weak compared to the standard deviation of the frequency distribution, then the oscillators cannot synchronize and the QL states remain mixed. When the coupling is very large, then the oscillator system is locked in phase. In this regime, dynamics in the QL state space can be unitary. That is, when the oscillators remain synchronized, the dynamics preserve purity.
@@
In the section, we have generalized our prior model, allowing any phase for the edges. This enabled us to consider explicitly the extent of phase synchronization in the classical network, which was shown to control the mixedness of the emergent QL states. Hence, this property of the variables defined by the classical network influences an important property of the states—phase coherence of the basis states is essential for sustaining stable superposition states. Depending on the properties of the of the classical phase oscillator network, the QL states can exhibit purity-preserving dynamics, or they can be completely mixed and undergo classical dynamics, or any case in between. When the oscillator network is synchronized, that is, the phases are coherent, then the QL states have high purity.
V.5 Decoherence
@@
An obvious example is a network of uncoupled oscillators with a frequency distribution. Starting in-phase, free evolution of the oscillators will dephase the order parameter and states in the network. The mechanism for such decoherence is similar to that for decoherence of a qubit[22]. Each coefficient in the chosen measurement basis of the effective state acquires a phase offset from the mean as a function of time. Let’s write this as $c_{m}^{i}=e^{i\zeta_{m}t}$. Then an off-diagonal density matrix element in this basis evolves as
@@
In the average, the phases in the emergent states evolve randomly because of the complexity of the underlying classical system, which leads to loss of phase coherence in these off-diagonal density matrix elements (they decay). This decoherence is independent of the measurement basis, unless we set the system up so that there is a preferred basis.
@@
The model captures the concept of an isolated QL system that succumbs to decoherence spontaneously. The numerical study shows an example of an initial pure state that decays to a final, equilibrium, state that is fully mixed in the product basis. The superpositions between basis states are completely lost, as evidenced by decay of the off-diagonal elements of the density matrix in the product basis, Fig. 5c. The classical variables—the phase oscillator network—control the coherence of the state, such that the quantum state space is slaved to the dynamics of the network. When the network is poorly synchronized, its many degrees of freedom act as a kind of environment that decoheres the states[22]. Imagine if we could only observe the state space, because the classical system was hidden somehow. Then we would observe spontaneous collapse of superpositions, Fig. 5b,c.
V.6 QL systems versus quantum systems
Here we speculate on what these studies of QL systems might inspire in terms of probing open questions about quantum systems.
Hidden variable theories in quantum mechanics aim to enable a physical interpretation of quantum mechanics, in particular to address questions such as how measurement works, what is the relationship between quantum and classical theories, and so on[54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. The main scope of hidden variable theories is to ask whether there is a deterministic system with predetermined outcomes prescribed for its properties, but this picture is hidden to us by the ways we measure the state vector. The motivation is that the hidden picture would clarify some of the mysteries that underpin the quantum theory. The no-hidden-variables theorems (e.g. Bell’s theorem) show how this is not the case[55].
The QL model gives a framework for examining hidden variables models in a new way. The reader will immediately object because Bell’s theorem, the Kochen-Specker theorem, and so on, are theorems that prove hidden variables cannot explain the perplexing outcomes of measurements on quantum systems. But, this result is consistent with how we find the interplay of the classical system and the QL states in the present work. It appears that correlations exhibited by the QL states are also built into the classical network. While the connections are preliminary at this stage, it does seem to be a topic worthy of additional scrutiny.
The main point is that as long as the classical system is constructed according to the same rules governing the construction of the quantum states—which is ensured by the QL graph—then the classical system and quantum states are bound such that we cannot arbitrarily assert maps on the QL states. That is, even if hidden variables do underlie a quantum system, they do not allow us to circumvent or rationalize observations that depend on quantum correlations. As an example, let’s consider the well-known no-cloning theorem[68].
@@ V.6 QL systems versus quantum systems
Recall that if $|a\rangle_{A}$ is a state of the QL graph $G_{A}$ and $|b\rangle_{B}$ is a state of the QL graph $G_{B}$, then $|a\rangle_{A}\otimes|b\rangle_{B}$ is a state of the Cartesian product of those graphs, $G_{A}\Box G_{B}$. The adjacency matrix of $G_{A}\Box G_{B}$ can be drawn in block form in such a way that the inherited edge structure from $G_{B}$ is highlighted on the diagonal:
$A_{A\Box B}=\begin{bmatrix}A_{B}&*&*&*\\ *&A_{B}&*&*\\ *&*&A_{B}&*\\ *&*&*&A_{B}\end{bmatrix}$ (36)
and all the other blocks, written as []delimited-[][*][ ∗ ] contain edge structure inherited from $G_{A}$.
When translated to graphs of QL states, the no-cloning theorem says that there is no unitary matrix that can transform $A_{A\Box B}$ such that, to an overall phase factor, diagonal blocks $[A_{B}]$ are mapped to copies of $[A_{A}]$, while leaving all the blocks []delimited-[][*][ ∗ ] unchanged.
@@
Now let’s consider the first []delimited-[][*][ ∗ ] block, which we require to be unchanged by the map because $|a\rangle_{A}$ should not be changed during the cloning except by a constant phase. The first two elements of the top row of []delimited-[][*][ ∗ ] are $e^{i(\theta_{m}-\theta_{1})}$ and $e^{i(\theta_{n}-\theta_{1})}$. Thus we require that $\theta_{m}-\theta_{1}=\theta_{n}-\theta_{1}$, that is, $\theta_{m}=\theta_{n}$. Clearly this us impossible to achieve if we want the $[A_{B}]$ blocks to be transformed. Therefore, even though the underlying network encoded by the graph is classical, there is no way of changing the phases of its constituent phase oscillators that allows a QL bit state to be cloned.
Another program in quantum theory is to understand how superpositions might collapse under dynamics governed by a modified version of the Schrödinger equation[69, 70, 71, 3]. Here the challenge is to understand how superpositions are lost, which is impossible under the linear evolution of the Schrödinger equation. Solving that problem would explain why macroscopic systems are not found in superposition and the theory would thus demystify the quantum to classical transition. One proposal is that the Schrödinger equation is modified so that at the micro-scale it appears unchanged from our usual formulation, but at longer length scales a nonlinear, stochastic feature emerges[70]. This new feature acts to collapse superpositions. Details of the nonlinear feature are hidden when we measure the state vectors.
The physical mechanisms for dynamical evolution of QL states suggests a connection might be made to this topic, in particular, it might inspire hypotheses to explain a mechanism for how superposition states can spontaneously collapse[3, 71, 70, 72]. Commonly, researchers incorporate a term into the Schrödinger equation that has no effect for typical quantum states, but for distant superpositions a stochastic, nonlinear term emerges that can initiate collapse of superpositions. The QL model considered in the present work appears to produce a similar result. We posited a somewhat different mechanism than proposed in prior important papers[3, 71, 70, 72] because in the present work the nonlinearity comes from the dynamics of the explicit classical system underlying the state space and it only controls phases of the underlying variables. The crucial parameter that impacts the states is the phase coherence of the classical network. However, the net result would be a modification of the Schrödinger equation prediction of dynamics in the state space. Indeed, we have noted how, when the oscillator network is poorly synchronized, it acts as a kind of bath for the state space. Thereby, it introduces a mechanism for decoherence similar to models for open quantum systems[73], where a system interacts with an external bath comprising many degrees of freedom.
V.7 Discussion
@@
The underlying classical system not only provides a mechanism for producing the QL state space, but, in addition, it actively controls the states. Owing to the complexity of the classical system (the system of coupled phase oscillators), it serves as a kind of bath, endowing the QL system with intrinsic properties like those characteristic of an open quantum system. For example, the system may exhibit intrinsic decoherence[74]. In this moderate coupling regime for the phase oscillator system, that is, the choice of coupling, $K$, relative to the distribution of natural frequencies of the oscillators, the state space undergoes dynamics that we call QL dynamics. The states are actively controlled by the dynamics of underlying system, which might be hidden from an observer. The example shown in Fig. 4 is that the mixedness of the emergent QL state decreases as the oscillators in the classical system synchronize. Whereas, in Fig. 5 we show a case where the mixedness of the emergent QL state increases as the oscillators in the classical system de-synchronize.
The model studied here suggests a way to understand a smooth transition in the state space from QL entangled states to mixed states. This would be observed if the propensity of the classical network to synchronize diminishes as the size of the system increases. An open question, therefore, is whether (and how) there can be a distance dependence governing the decoherence so that distant superpositions naturally collapse. For example, we can imagine that the noise that gives the oscillators the frequency distribution is correlated on molecular length scales, but uncorrelated on longer length scales. This setup might serve to collapse distant superpositions, but preserve superpositions among closely-spaced systems.
@@
Another possibility is that synchronization diminishes, after some cut-off, as more and more QL graphs are included in the product graph. A setup with this property (to consider in future work) is to endow the network and corresponding graph with a cut-off distance for couplings, or edges. In other words, the graph is $d$-regular, but only over a certain spatial extent and for the vertices that are included in that spatial area. The graph is everywhere $d$-regular, but two domains that are separated by a distance much greater than the cutoff are not fully connected. We can imagine this as a window that we can slide over the graph such that anywhere within the window the graph is $d-$regular. In this scenario, synchronization has a size limit. So, when the network enlarges as more QL bits are included in the product, there will be a point where the entire state space cannot be coherent. The poorly connected network domain will limit the QL state space by acting as a bath and introducing decoherence. This kind of construction might give a small system—that is, smaller than the ‘window’—that is QL, but as the system is enlarged, the properties of its state space would become more classical (mixed). Here we have in mind that synchronizability of the classical network is a function of the number of vertices.
VI Conclusions

### v4

@@ V.1 Phase and the adjacency matrix
This unitary transformation of $A$ does not change the spectrum, but it does rotate the basis for the eigenstates. We use this transformation below, after defining a projection of the emergent states, to show how the eigenvectors of the projected emergent states transform under continuous rotations. Later in this paper, we use the transformation of Eq. 13 to account for how oscillator phases in the classical network dictate the eigenvector basis. That basis change, in general, varies throughout the ensemble and therefore influences the mixedness of the state[42]. We use the term ‘ensemble’ to mean an average over initial phases to account for state preparation.
@@
In the numerical examples shown below we obtain the phases from simulations using the Kuramoto model, described in the previous section. For any time point during the simulation, we use the oscillator phases to define $\Phi$, then we diagonalize $\Phi^{-1}A\Phi$ to obtain the eigenvectors of the emergent states—now with complex coefficients. We project these coefficients to those for an effective $2^{q}$-state system in the product basis for $q$ QL bits. As a starting point, we now describe for the example of a single QL bit. The method is generalized in the obvious way for a system comprising any number of QL bits, as shown in Sec. V.C.
V.2 The effective 2-state model for QL bit emergent states
@@ V.3 Phase in the states
For a synchronized oscillator system on the QL bit graph with all edges set to 1 in the adjacency matrix, we find that $\alpha_{0}=\beta_{0}\approx 1/\sqrt{2}$ and $\alpha_{1}=-\beta_{1}\approx 1/\sqrt{2}$ (we use the convention that $\beta_{1}$ takes the negative sign). More generally, we can bias the edges that connect the two subgraphs by giving them values (in the adjacency matrix) of any value on the unit circle in the complex plane. Recall that the special case of negative-signed edges connecting the subgraphs can be interpreted as phases of the two subgraphs being perfectly out of phase[1]. The resulting eigenvectors can be evaluated by considering the effective adjacency matrix for the QL bit graph comprising two $d$-regular subgraphs coupled by edges given the bias $e^{i\gamma}$:
@@
By rotating the edge bias, we therefore generate states (when the system is perfectly synchronized) that emulate, for example, the states of a system of spins or polarization states of photons. The notable emergent states isomorphic to polarization states of light are: (i) the linear combinations $a_{1}\pm a_{2}$, when edges take values ±1$\pm$ 1± 1 respectively, which are comparable to linear combinations of vertical and horizontal polarization; (ii) the states $a_{1}\pm ia_{2}$, when edges take values of ±$\pm i$, analogous to right- and left-circularly polarized light.
V.4 Synchronization of quantum bits
@@
with the $\mathbf{L}$ being the $\times 4$ 4 skew symmetric matrices[47]:
@@
VI Synchronization versus dephasing of QL states
VI.1 Methods
Here we study a phase oscillator network model with a coupling structure templated by the Cartesian product of two QL bit graphs $G_{A}\Box G_{B}$. The phase oscillators are indicated by vertices of the graph, and are assigned a phase that biases the edges.
@@ VI.1 Methods
VI.2 Parameters and State preparation
@@
In sum, whether or not a system of oscillators will perfectly synchronize depends not only on $K$ and $\sigma_{\nu}$, but also on the initial conditions. That is, the specific assignment of frequencies and initial phases to each oscillator in the network. For insight into this statement see, for example, Refs. [9, 10, 8, 49, 50]. This sensitivity of the system to initial conditions emphasizes the nonlinearity of the network.
VI.3 States of a two-QL bit composite system
To provide context for the QL state we study in the remainder of the paper, let’s recall the following orthonormal basis for a composite system in $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and equate it to the symbols we use for the analogous QL state basis:
$|0_{A}\rangle\otimes|0_{B}\rangle=|a_{1}\rangle|b_{1}\rangle$
$|1_{A}\rangle\otimes|0_{B}\rangle=|a_{2}\rangle|b_{1}\rangle$
$|0_{A}\rangle\otimes|1_{B}\rangle=|a_{1}\rangle|b_{2}\rangle$
$|1_{A}\rangle\otimes|1_{B}\rangle=|a_{2}\rangle|b_{2}\rangle.$
Any state $|\Psi\rangle$ in $\mathcal{H}$ can be written as a linear combination of these basis states:
$|\Psi\rangle=c_{1}|0_{A}\rangle\otimes|0_{B}\rangle+c_{2}|1_{A}\rangle\otimes|0_{B}\rangle+c_{3}|0_{A}\rangle\otimes|1_{B}\rangle+c_{4}|1_{A}\rangle\otimes|1_{B}\rangle.$ (25)
Furthermore, if $|\Psi\rangle$ is a separable (pure) state, then we can decompose it as $|\Psi\rangle=|\psi_{A}\rangle\otimes|\psi_{B}\rangle$, where $|\psi_{A}\rangle=\alpha_{A}|0_{A}\rangle+\beta_{A}|1_{A}\rangle$ and $|\psi_{B}\rangle=\alpha_{B}|0_{B}\rangle+\beta_{B}|1_{B}\rangle$. That is, these are the same as the effective states of the QL bit, $W_{i}^{2\times 2}$, Eq. 19. The separable states of $|\Psi\rangle$ indexed by $i$, written below as $W_{i}^{4\times 4}$, are generated by the Cartesian product of two QL bit graphs. As introduced earlier in the paper, here we focus on just one of the $W_{i}^{4\times 4}$ states, that with largest eigenvalue. It corresponds to the ‘ground state’ of the classical phase oscillator system (according to the way we write Eq. 10), and gives the instantaneous eigenvector of the dynamical network projected to the effective four-state model. This eigenvector does not vary with time once the phase oscillator system reaches equilibrium.
To obtain this instantaneous eigenvector of the dynamical network projected to the effective four-state model, we take the oscillator phases and construct the unitary matrix $\Phi$, then transform the adjacency matrix of the graph according to Eq. 12. Then we project this emergent state with highest eigenvalue onto the basis where each subgraph, labeled $a_{1}$, $a_{2}$, etc., maps to the corresponding single state $|a_{1}\rangle$, $|a_{2}\rangle$, etc. That gives all four of the $W_{i}^{4\times 4}$, in terms of the complex coefficients $c^{i}_{k}(t)$ obtained from the projections. The density matrix for the emergent state is calculated in this basis by statistically averaging over the pure states:
$W_{i}^{4\times 4}(t)=c^{i}_{1}(t)|a_{1}\rangle|b_{1}\rangle+c^{i}_{2}(t)|a_{2}\rangle|b_{1}\rangle+c^{i}_{3}(t)|a_{1}\rangle|b_{2}\rangle+c^{i}_{4}(t)|a_{2}\rangle|b_{2}\rangle,$ (26)
@@ VI.3 States of a two-QL bit composite system
for each time point $t$ in a simulation. We obtain the coefficients $c^{i}_{k}(t)$ as follows. We form the Cartesian product of two QL bit graphs, say $G_{A}$ and $G_{B}$, each comprising two subgraphs with $n_{0}$ vertices. From the product, $G_{A}\Box G_{B}$, we end up with four subgraphs (each with $n_{0}^{2}$ vertices). The subgraphs are connected with a certain structure of edges that define the correlations. The vertices in each subgraph are labeled by all possible $n_{0}^{2}$ pairs of vertex indices, where one index is taken from a subgraph of $G_{A}$ and the other from a subgraph of $G_{B}$. We project the emergent states of the entire graph to effective states in a chosen basis. Here we choose the natural product basis associated directly with the subgraphs of $G_{A}\Box G_{B}$: $|a_{1}\rangle|b_{1}\rangle$, $|a_{2}\rangle|b_{1}\rangle$, $|a_{1}\rangle|b_{2}\rangle$, $|a_{2}\rangle|b_{2}\rangle$. That basis is selected by defining the vectors
@@
where $\sum_{\text{prep}}$ means an average over state preparations. We report $\mathrm{Tr}(\rho)^{2}$ as a simple measure of mixedness of the state. Notice that the complex phases cancel to unity in the diagonal elements of $\rho_{mn}(t)$. On the other hand, the complex phases are retained in the off-diagonal elements and unless the system is synchronized, they interfere in the average and can thereby diminish or erase the off-diagonal elements, introducing state mixedness[42].
The vectors $J_{a1b1}$ and so on define the basis for the QL state with respect to the subgraphs. Here we set them up so that the basis is the product basis associated directly with the subgraphs of $G_{A}\Box G_{B}$. In the following sections we analyze these states theoretically. We do not impose any kind of measurement formalism in the present work because we simply aim to study the properties of these states as a function of the properties of the classical phase oscillator network.
VI.4 Synchronization
@@
Despite the fact that the underlying synchronization dynamics are inherently nonlinear, maps in the state space taking the state from one time point to another, $\Xi:\rho(t_{1})\rightarrow\rho(t_{2})$, are linear maps. As discussed earlier in the paper, the reason for this linear representation of a nonlinear systems boils down to the way Eq. 12 evolves the adge biases in the adjacency matrix by a unitary operator.
@@
The dynamical evolution is generally not unitary; being unitary only in the limit of strong synchronization. Therefore, different regimes can be identified. When all the oscillators are strongly synchronized within each QL bit, but not necessarily between QL bits, then the general model for the single-excitation subspace specializes to the Lohe model. Single-excitation subspace means that the network graph carries the structure of the tensor sum of QL bit graphs and QL bits are coupled to other QL bits. When the coupling $K$ for the classical system is very weak compared to the standard deviation of the frequency distribution, then the oscillators cannot synchronize and the QL states remain mixed. When the coupling is very large, then the oscillator system is locked in phase. In this regime, dynamics in the QL state space can be unitary. That is, when the oscillators remain synchronized, the dynamics preserve purity.
@@
In this section, we have generalized our prior model, allowing any phase for the edges. This enabled us to consider explicitly the extent of phase synchronization in the classical network, which was shown to control the mixedness of the emergent QL states. Hence, this property of the variables defined by the classical network influences an important property of the states—phase coherence of the basis states is essential for sustaining stable superposition states. Depending on the properties of the classical phase oscillator network, the QL states can exhibit purity-preserving dynamics, or they can be completely mixed and undergo classical dynamics, or any case in between. When the oscillator network is synchronized, that is, the phases are coherent, then the QL states have high purity.
VI.5 Dephasing
@@
An obvious example is a network of uncoupled oscillators with a frequency distribution. Starting in-phase, free evolution of the oscillators will dephase the order parameter and states in the network. The mechanism for such dephasing is similar to that for decoherence of a qubit[52]. Each coefficient in the chosen basis for the effective states acquires a phase offset from the mean as a function of time. Let’s write this as $c_{m}^{i}=e^{i\zeta_{m}t}$. Then an off-diagonal density matrix element in this basis evolves as
@@
In the average, the phases in the emergent states evolve randomly because of the complexity of the underlying classical system, which leads to loss of phase coherence in these off-diagonal density matrix elements (they decay). This dephasing is independent of how we project the effective states.
@@
The model captures the concept of an isolated QL system that succumbs to dephasing spontaneously. The numerical study shows an example of an initial pure state that decays to a final, equilibrium, state that is fully mixed in the product basis. The superpositions between basis states are completely lost, as evidenced by decay of the off-diagonal elements of the density matrix in the product basis, Fig. 6c. The classical variables—the phase oscillator network—control the coherence of the state, such that the quantum state space is slaved to the dynamics of the network. When the network is poorly synchronized, its many degrees of freedom act as a kind of environment that decoheres the states[52]. Imagine if we could only observe the state space, because the classical system was hidden somehow. Then we would observe spontaneous collapse of superpositions, Fig. 6b,c.
VI.6 Discussion
We found that, for the particular example of the phase oscillator network, that nonlinear dynamics of the classical system are represented by linear dynamics in the space of QL states. That result comes from the way the dynamics of the phase oscillators are represented as unitary maps on the adjacency matrix that generates the states. It provides a useful example of how the QL states can be constructed so they mimic quantum states rather closely. As another example, let’s consider the well-known no-cloning theorem[53].
@@ VI.6 Discussion
Recall that if $|a\rangle_{A}$ is a state of the QL graph $G_{A}$ and $|b\rangle_{B}$ is a state of the QL graph $G_{B}$, then $|a\rangle_{A}\otimes|b\rangle_{B}$ is a state of the Cartesian product of those graphs, $G_{A}\Box G_{B}$. Recall that each vertex of $G_{A}$ (alternatively $G_{B}$) is now the graph $G_{B}$ (alternatively $G_{A}$), with additional edges added appropriately. The adjacency matrix of $G_{A}\Box G_{B}$ can be drawn in block form in such a way that the inherited edge structure from $G_{B}$ is highlighted on the diagonal:
$A_{A\Box B}=\begin{bmatrix}A_{B}&*&*&*&\dots\\ *&A_{B}&*&*&\dots\\ *&*&A_{B}&*&\dots\\ *&*&*&A_{B}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (35)
and all the other blocks, written as [][*][ ∗ ] contain edge structure inherited from $G_{A}$.
When translated to graphs of QL states, the no-cloning theorem says that there is no unitary matrix that can transform $A_{A\Box B}$ such that, to an overall phase factor, diagonal blocks $[A_{B}]$ are mapped to copies of $[A_{A}]$, while leaving all the blocks [][*][ ∗ ] unchanged.
@@
Now let’s consider the first [][*][ ∗ ] block, which we require to be unchanged by the map because $|a\rangle_{A}$ should not be changed during the cloning except by a constant phase. The first two elements of the top row of [][*][ ∗ ] are $e^{i(\theta_{m}-\theta_{1})}$ and $e^{i(\theta_{n}-\theta_{1})}$. Thus we require that $\theta_{m}-\theta_{1}=\theta_{n}-\theta_{1}$, that is, $\theta_{m}=\theta_{n}$. Clearly this us impossible to achieve if we want the $[A_{B}]$ blocks to be transformed. Therefore, even though the underlying network encoded by the graph is classical, there is no way of changing the phases of its constituent phase oscillators that allows a QL bit state to be cloned.
This result is not meant to be profound, but its significance is that it illustrates how the structure of the adjacency matrix of a composite QL system is the important factor that produces effective states that are ‘quantum like’. The structure is inherited by the special classical network that is identified by the graph representation. In that regard, here we constructed the classical network by an easy one-to-one map from the graph, where each vertex becomes an oscillator and edges show which oscillators are coupled to which other oscillators. Note though, that the classical system is very large—it has many vertices—but we map it to a much small state space because the number of QL bits that produce the composite quantum system is much fewer than the number of vertices in the graph. We are free to represent the graph in many other ways in the classical system. The classical system could be even larger than the graph, or it could be given in terms of continuous variables rather than a finite number of discrete vertices.
@@
In the present study of dynamics of the network, we found that the underlying classical system not only provides a mechanism for producing the QL state space, but, in addition, it actively controls the states. Owing to the complexity of the classical system (the system of coupled phase oscillators), we could identify different kinds of evolution of the QL states. In a particular regime where the oscillators do not synchronize, because the couplings $K$ between the classical phase oscillators are small, the phase oscillator network serves as a kind of bath, endowing the QL system with intrinsic properties like those characteristic of an open quantum system. For example, the system may exhibit intrinsic decoherence[54]. That is seen in Fig. 6, where the mixedness of the emergent QL state increases as the oscillators in the classical system de-synchronize. In contrast, when $K$ is large enough, the classical network spontaneously synchronizes. Them as shown in Fig. 5, the mixedness of the emergent QL state decreases as the oscillators in the classical system synchronize (the purity of the state increases).
@@
The model studied here suggests a way to understand a smooth transition in the state space from QL entangled states to mixed states. This would be observed if the propensity of the classical network to synchronize diminishes as the size of the system increases. An open question, therefore, is whether (and how) there can be a distance dependence governing the decoherence so that distant superpositions naturally collapse. For example, we can imagine that the noise that gives the oscillators the frequency distribution is correlated on short length scales relative to the distance between oscillators, but uncorrelated on longer length scales. This setup might serve to collapse distant superpositions, but preserve superpositions among closely-spaced systems. Another possibility is that synchronization diminishes, after some cut-off, as more and more QL graphs are included in the product graph. A setup with this property (to consider in future work) is to endow the network and corresponding graph with a cut-off distance for couplings, or edges. In other words, the graph is $d$-regular, but only over a certain spatial extent and for the vertices that are included in that spatial area. The graph is everywhere $d$-regular, but two domains that are separated by a distance much greater than the cutoff are not fully connected. We can imagine this as a window that we can slide over the graph such that anywhere within the window the graph is $d-$regular. In this scenario, synchronization has a size limit. So, when the network enlarges as more QL bits are included in the product, there will be a point where the entire state space cannot be coherent. The poorly connected network domain will limit the QL state space by acting as a bath and introducing decoherence. This kind of construction might give a small system—that is, smaller than the ‘window’—that is QL, but as the system is enlarged, the properties of its state space would become more classical (mixed). Here we have in mind that synchronizability of the classical network is a function of the number of vertices.
VII Conclusions